OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

@jim_thompson5910 @Hero @mathstudent55 help please?

hero (hero):

Start with f(3) and work our way up

OpenStudy (anonymous):

F(3) would be 4? Or is it saying to substitute n with 5? I'm completely lost on how to do this.

hero (hero):

Since f(1) = 2 and f(2) = 3, substitute those in to the given equation f(n) = f(1) + f(2) + f(n - 1) to get f(n) = 2 + 3 + f(n - 1) which simplifies to f(n) = 5 + f(n - 1) So far f(3) = 5 + f(3 - 1) f(3) = 5 + f(2) f(3) = 5 + 3 f(3) = 8

OpenStudy (anonymous):

So 4 would be 13?